# Basis Construction for Spline Spaces over Arbitrary Partitions from a Dimensional Stable Perspective

### v1

@@
Abstract. This paper introduces a novel framework for constructing $C^{r}$ basis functions for polynomial spline spaces of degree $d$ over arbitrary planar polygonal partitions, overturning the belief that basis functions cannot be constructed on dimensionally unstable meshes. We provide a comprehensive comparison of basis construction methods, classifying them as explicit, semi-explicit, and implicit. Our method, a semi-explicit construction using Extended Edge Elimination conditions, uniquely resolves all theoretical challenges in spline spaces by ensuring a complete basis. For the first time, we construct basis functions for the spline space over the Morgan-Scott partition, previously unachieved, and elucidate dimensional instability through this construction.
@@ 1 Introduction
Polynomial spline spaces defined over planar polygonal partitions, such as T-meshes and triangulations, have been a subject of study since the 1970s (Strang, 1973; Schumaker, 1979; Nürnberger and Zeilfelder, 2000). These spaces are significant in approximation theory, numerical analysis (Lai and Schumaker, 2007; Schumaker, 2007), and isogeometric analysis (Cottrell et al., 2009), particularly in geometric modeling (Cohen et al., 2001) and engineering applications (Cottrell et al., 2009). Research on polynomial spline spaces typically focuses on two fundamental problems: dimension calculation and basis construction.
@@
For dimension calculation, spline spaces over triangulations have been extensively studied, with key results from the 1970s to 2000s summarized in Section 3 of Nürnberger and Zeilfelder (2000) and recent advancements in Section 2 of Grošelj and Speleers (2023); Lyche et al. (2025). For T-meshes, significant progress has been made over the past two decades (Deng et al., 2006; Li et al., 2006; Mourrain, 2014; Li and Deng, 2016). However, a critical challenge in both triangulation and T-mesh spline spaces is dimensional instability (Morgan and Scott, 1977; Alfeld, 1986; Gmelig Meyling and Pfluger, 1985; Shi, 1991; Deng et al., 2000; Li and Chen, 2011; Berdinsky et al., 2012; Huang and Chen, 2024, 2025). This instability implies that the dimension of highly smooth spline spaces depends not only on topological properties but also on geometric configurations of the polygonal partition. A notable example is the Morgan-Scott triangulation (Morgan and Scott, 1977).
@@
In basis construction, dimensional instability in highly smooth spline spaces over specific partitions complicates the direct construction of basis functions. Traditionally, researchers have sought to avoid issues arising from dimensional instability when constructing basis functions. For spline spaces over general polygonal partitions, basis functions have been constructed for cross-cut and quasi-cross-cut partitions (Wang, 2013; Manni, 1992). Recent advancements for triangulations include new constructions (Speleers, 2013; Lyche et al., 2022; Grošelj and Speleers, 2023, 2017; Lyche et al., 2025). For T-meshes, notable constructions include T-splines (Sederberg et al., 2003, 2004; Li et al., 2012), HB-splines (Forsey and Bartels, 1988; Vuong et al., 2011), LR-splines (Dokken et al., 2013), and PHT-splines (Deng et al., 2008). However, except for PHT-splines, these typically do not form a complete basis for the polynomial spline space over the corresponding T-mesh (Buffa et al., 2010; Mokriš et al., 2014; Giannelli and Jüttler, 2013; Dokken et al., 2013).
@@
This paper extends the framework of to polynomial spline spaces over arbitrary planar polygonal partitions. We observe that any polygonal partition can be transformed into a cross-cut or quasi-cross-cut partition, which possess stable dimensions and well-defined basis function structures (Wang, 2013; Manni, 1992). By applying the Extended Edge Elimination Condition (EEE condition), we derive a basis for the spline space over the original partition. To address the complexity of the EEE condition due to excessive extensions, we propose a simplified framework based on dimensional stability. This framework also facilitates the analysis of dimensional instability.
@@
This study advances basis construction for spline spaces over arbitrary polygonal partitions, challenging the notion that basis functions cannot be constructed on dimensionally unstable meshes. The key contributions are:
@@ 2 Preliminaries
Lemma 2.1
(Chui and Wang, 1983a) Let $\Delta_{c}$ be a cross-cut partition of $\Omega$ with $L$ cross-cuts and $V$ interior vertices $A_{1},\ldots,A_{V}$ in $\Omega$ such that $N_{i}$ cross-cuts intersect at $A_{i}$ for $i=1,2,\ldots,V$. Then the dimension of spline space over $\Delta_{c}$ is
@@
$k_{d}^{\mu}(N_{i})=\sum\limits_{j=1}^{d-\mu}\left(N_{i}(d-\mu-j+1)-(d-j+2)\right)_{+}.$ (2.1)
Lemma 2.2
(Chui and Wang, 1983a; Manni, 1992) Let $\Delta_{qc}$ be a quasi-cross-cut partition of $\Omega$ with $L_{1}$ cross-cuts, $L_{2}$ rays and $V$ interior vertices $A_{1},\ldots,A_{V}$ in $\Omega$ such that $N_{i}$ cross-cuts and rays intersect at $A_{i}$ for $i=1,2,\ldots,V$. Then the dimension of spline space over $\Delta_{qc}$ is
@@
Lemma 2.3
(Wang, 2013) For any polygonal partition $\Delta$,let $\phi_{i},\phi_{j}$ are two adjcent cells of $\Delta$, $s(x,y)|_{\phi_{i}}=s_{i}(x,y),s(x,y)|_{\phi_{j}}=s_{j}(x,y)$, and the equation of the common edge of $\phi_{i}$ and $\phi_{j}$ is $L_{ij}(x,y)=0$. Then
@@ 3 Framework for Basis Function Construction
In this section, we introduce the concept of a Construction-Suitable Partition for spline spaces, enabling direct basis construction without edge extension or elimination. We apply the Extended Edge Elimination (EEE) conditions to this polygonal partition, establishing a framework for constructing basis functions for the spline space. To clarify the construction process, we first consider the one-dimensional case before addressing the two-dimensional case.
@@ 3.1 Motivation: One-Dimensional Case
Definition 3.1
(Lyche and Morken, 2008) A knot vector $\Delta_{1}$ is a refinement of $\Delta_{2}$ if every real number in $\Delta_{2}$ appears in $\Delta_{1}$ with at least the same multiplicity.
@@
This study aims to construct a basis for the spline space $S_{d}^{k}(\Delta_{2})$ from $S_{d}^{k}(\Delta_{1})$, addressing the following problem:
$\textbf{Construct a basis for }S_{d}^{k}(\Delta_{2})\textbf{ from }S_{d}^{k}(\Delta_{1}).$ (3.2)
A key observation is that the primary distinction between $S_{d}^{k}(\Delta_{1})$ and $S_{d}^{k}(\Delta_{2})$ lies in the smoothness condition: in $S_{d}^{k}(\Delta_{2})$, the knots in $\Delta_{1}\setminus\Delta_{2}$ are $C^{\infty}$. This leads to the following lemma for $S_{d}^{k}(\Delta_{2})$.
Lemma 3.2
@@
$\left.\frac{\partial^{d}_{+}}{\partial x^{d}}s(x)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}=\left.\frac{\partial^{d}_{-}}{\partial x^{d}}s(x)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}.$ (3.3)
@@
$\left.\sum c_{i}\left(\frac{\partial^{d}_{+}}{\partial x^{d}}B_{i,\Delta_{1}}(x)-\frac{\partial^{d}_{-}}{\partial x^{d}}B_{i,\Delta_{1}}(x)\right)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}=0,$ (3.4)
@@
$\left\{\sum c_{i,j}B_{i,\Delta_{1}}(x)\right\}_{j=1}^{h}$ (3.5)
@@ 3.2 Two-Dimensional Case
In (Zhong et al., 2025), a basis construction method for spline spaces over diagonalizable T-meshes is introduced, followed by the use of EEE conditions to construct bases for spline spaces over arbitrary T-meshes. This section generalizes the basis construction for diagonalizable T-meshes, introducing the concept of Construction-Suitable Partition. We then clarify the relationship and distinctions between Dimensional-Stable Partitions and Construction-Suitable Partitions. Finally, we extend the EEE conditions to arbitrary polygonal partitions, enabling basis construction for spline spaces over such partitions. Before introducing the main result, we examine the extended polygonal partition, denoted $\mathrm{ext}_{s}(\Delta)$, as follows.
Definition 3.2
@@
$\mathrm{ext}_{0}({\Delta}):={\Delta}\subseteq\mathrm{ext}_{1}({\Delta})\subseteq\mathrm{ext}_{2}({\Delta})\subseteq\cdots\subseteq\mathrm{ext}_{t}({\Delta}):={\Delta}_{c}.$
@@
$S_{d}^{\mu}\left(\mathrm{ext}_{0}({\Delta})\right)\subseteq S_{d}^{\mu}\left(\mathrm{ext}_{1}({\Delta})\right)\subseteq\cdots\subseteq S_{d}^{\mu}\left(\mathrm{ext}_{t}({\Delta})\right).$
where $\subseteq$” denotes a subspace relation.
We now define the Construction-Suitable Partition (or Construction-Suitable Mesh), which enables direct construction of basis functions for the spline space $S_{d}^{\mu}({\Delta})$ and serves as a core component in the basis construction framework.
Definition 3.3
A polygonal partition $\Delta$ is termed a Construction-Suitable Partition if:
@@
By Definitions 3.3 and 3.4, every Construction-Suitable Partition is a Dimensional-Stable Partition, but the converse is not true. The Construction-Suitable Partition depends on the basis construction method. A Dimensional-Stable Partition may form a Construction-Suitable Partition under one construction method but not under another. Whatever, the dimensional stability condition is necessary.
Example 3.1
@@
(Zhong et al., 2025) Consider the T-mesh $\Delta$} with a T-cycle, as shown in the left subfigure of Figure 1. For the spline space $S_{2,2}^{1,1}({\Delta})$, the dimension is stable. We construct basis functions for $S_{2,2}^{1,1}({\Delta})$ using two methods:
@@
  • U-spline method (Thomas et al., 2022): Basis functions are constructed by forming linear systems from smoothness constraints on the Bézier coordinates of each cell and solving these systems.
@@
    For the T-mesh $\Delta$} depicted in the left subfigure of Figure 1, this method directly yields a basis for $S_{2,2}^{1,1}({\Delta})$, confirming that $\Delta$} is a Construction-Suitable Partition.
@@
  • PT-spline method (Zhong et al., 2025): Basis functions are constructed by extending edges in the T-mesh to form an extended T-mesh $\mathrm{ext}_{s}({\Delta})$ that supports local tensor product B-splines, satisfying the dimension formula for $S_{\boldsymbol{d}}^{\boldsymbol{\mu}}(\mathrm{ext}_{s}({\Delta}))$. The Extended Edge Elimination Condition is then applied to obtain a basis for $S_{\boldsymbol{d}}^{\boldsymbol{\mu}}(\mathrm{ext}_{s}({\Delta}))$.
@@
    For the T-mesh $\Delta$} depicted in the left subfigure of Figure 1, edge extensions are necessary to satisfy the PT-spline construction conditions, as depicted in the right subfigure of Figure 1. Consequently, the extended T-mesh $\mathrm{ext}_{2}({\Delta})$ forms a Construction-Suitable Partition. See Example 11 in (Zhong et al., 2025) for details.
Dimensional-Stable PartitionConstruction-Suitable Partition
@@
Figure 1: Comparison of Dimensional-Stable Partition and Construction-Suitable Partition
Definition 3.5
@@
For an arbitrary polygonal partition $\Delta$, let $\mathcal{M}$ denote a basis construction method. The pair $(\Delta,\mathcal{M})$ is compatible if $\Delta$ is a construction-suitable partition under $\mathcal{M}$.
@@
Many partitions studied for constructing spline basis functions are Construction-Suitable Partitions, enabling direct basis construction. In the T-mesh category, these include hierarchical meshes with mild restrictions for HB-splines Mokriš et al. (2014), AS T-meshes for AS T-splines Li et al. (2012); Li and Scott (2014), restricted LR-meshes for LR-splines Patrizi et al. (2020), and hierarchical T-meshes for PHT-splines Deng et al. (2008). In the triangulation category, examples include Clough-Tocher (CT) splits Lai and Schumaker (2007); Sablonnière (1987) for $S_{d}^{2}(\Delta_{\mathrm{CT}})$ with $d\geq 7$ and Powell-Sabin (PS) 6 and 12 splits Powell and Sabin (1977); Alfeld and Schumaker (2002); Schumaker and Sorokina (2006); Lai and Schumaker (2007); Speleers (2013) for $S_{d}^{2}(\Delta_{\mathrm{PS}})$ with $d\geq 5$ etc.
Example 3.2
@@
The pair $(\Delta,\mathcal{M}_{\text{HB}})$ is not compatible because $\Delta$ does not form a hierarchical B-spline mesh. For $S_{2}^{1}(\Delta)$, $(\Delta,\mathcal{M}_{\text{PHT}})$ is not compatible as the smoothness and degree do not satisfy PHT-spline requirements, which mandate smoothness less than half the degree. For $S_{3}^{1}(\Delta)$, $(\Delta,\mathcal{M}_{\text{PHT}})$ is compatible since $\Delta$ is a hierarchical T-mesh and the smoothness meets PHT-spline requirements.
@@
In fact, cross-cut and quasi-cross-cut partitions are both Construction-Suitable Partitions for the construction method mentioned in Wang (2013). The cross-cut partition is the maximal Construction-Suitable Partition for a given polygonal partition $\Delta$.
@@
Definition 3.6
Given a polygonal partition $\Delta$} and its extended Construction-Suitable Partition $\mathrm{ext}_{s}({\Delta})$, as defined in Definition 3.3, assume that $\{N_{1},N_{2},\ldots,N_{n}\}$ forms a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}({\Delta}))$. For an extended edge $l$ in $\mathrm{ext}_{s}({\Delta})$, let $\phi_{1}$ and $\phi_{2}$ be two adjacent cells intersecting at $l$. The condition
$\sum_{i=1}^{n}c_{i}N_{i}\in C^{d}(\overline{\phi_{1}\cup\phi_{2}})$ (3.6)
@@
$\sum_{i=1}^{n}c_{i}\cdot\frac{\partial^{r}}{\partial\boldsymbol{n}^{r}}\left(N_{i}|_{\phi_{1}}-N_{i}|_{\phi_{2}}\right)\Big{|}_{l}\equiv 0,\quad r=\mu+1,\ldots,d,$ (3.7)
@@
Proof  Since $N_{i}\in S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$, the EEE condition (3.6) implies
@@
Definition 3.7
Given a polygonal partition $\Delta$ and its extended construction-suitable partition $\mathrm{ext}_{s}(\Delta)$, let $\{N_{1},N_{2},\ldots,N_{n}\}$ be a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$. For the set of extended edges $\{l_{1},l_{2},\ldots,l_{s}\}$, the EEE conditions for each $l_{i}$, $i=1,\ldots,s$, form the linear system
$M\boldsymbol{c}=0,$ (3.8)
@@
Theorem 3.2
Given a polygonal partition $\Delta$ and its extended construction-suitable partition $\mathrm{ext}_{s}(\Delta)$, let $\{N_{1},N_{2},\ldots,N_{n}\}$ be a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$. If the homogeneous linear system $M\boldsymbol{c}=0$ from (3.8) has a basis of solutions $\{\tilde{c}_{j}\}_{j=1}^{h}$, where $\tilde{c}_{j}=(\tilde{c}_{1,j},\tilde{c}_{2,j},\ldots,\tilde{c}_{n,j})$, then the functions
@@
Proof  Let $\{l_{1},l_{2},\ldots,l_{s}\}$ denote the set of all extended edges. The proof proceeds in two steps:
@@
    For any $s\in S_{d}^{\mu}(\mathscr{T})$, Lemma 3.5 implies $s\in S_{d}^{\mu}(\mathrm{ext}_{s}(\mathscr{T}))$, allowing the representation
@@
    Since $s\in S_{d}^{\mu}(\mathscr{T})$, for each virtual edge $l_{i}$, $i=1,2,\ldots,s$, it follows $\sum_{i=1}^{n}c_{i}N_{i}\in C^{\infty}\left(\overline{\phi_{1}^{{(i)}}\cup\phi_{2}^{{(i)}}}\right)$ where $\phi_{1}^{{(i)}},\phi_{2}^{{(i)}}$ are two adjacent cells of $l_{i}$. Consequently, in $\mathscr{T}$, the following conditions hold:

### v2

@@
Abstract
This paper introduces a novel framework for constructing $C^{r}$ basis functions for polynomial spline spaces of degree $d$ over arbitrary planar polygonal partitions, overturning the belief that basis functions cannot be constructed on dimensional unstable meshes. We provide a comprehensive comparison of basis construction methods, classifying them as explicit, semi-implicit, and implicit. Our method, a semi-implicit construction using Extended Edge Elimination conditions, uniquely resolves fundamental theoretical challenges in the basis construction of spline spaces by ensuring a complete basis. For the first time, we construct explicit basis functions for the spline space over the Morgan-Scott partition, previously unachieved, and elucidate dimensional instability through this construction.
keywords:
Basis construction for spline space, Arbitrary partition, Dimensional stability, Isogeometric analysis.
{AMS}
65D17, 65D07, 65N30.
@@ 1 Introduction
Polynomial spline spaces defined over planar polygonal partitions, such as T-meshes and triangulations, have been a subject of study since the 1970s Schumaker (1979); Nürnberger and Zeilfelder (2000). These spaces are significant in approximation theory, numerical analysis Lai and Schumaker (2007); Schumaker (2007), and isogeometric analysis Cottrell et al. (2009), particularly in geometric modeling Cohen et al. (2001) and engineering applications Cottrell et al. (2009). Research on polynomial spline spaces typically focuses on two fundamental problems: dimension calculation and basis construction.
@@
For dimension calculation, spline spaces over triangulations have been extensively studied, with key results from the 1970s to 2000s summarized in Section 3 of Nürnberger and Zeilfelder (2000) and recent advancements in Section 2 of Grošelj and Speleers (2023); Lyche et al. (2025). For T-meshes, significant progress has been made over the past two decades Deng et al. (2006); Li et al. (2006); Mourrain (2014); Li and Deng (2016). However, a critical challenge in both triangulation and T-mesh spline spaces is dimensional instability Morgan and Scott (1977); Alfeld (1986); Gmelig Meyling and Pfluger (1985); Shi (1991); Deng et al. (2000); Li and Chen (2011); Berdinsky et al. (2012); Huang and Chen (2024, 2025). This instability implies that the dimension of highly smooth spline spaces depends not only on topological properties but also on geometric configurations of the polygonal partition. A notable example is the Morgan-Scott triangulation Morgan and Scott (1977).
@@
In basis construction, dimensional instability in highly smooth spline spaces over specific partitions complicates the direct construction of basis functions. Traditionally, researchers have sought to avoid issues arising from dimensional instability when constructing basis functions. For spline spaces over general polygonal partitions, basis functions have been constructed for cross-cut and quasi-cross-cut partitions Wang (2013); Manni (1992). Recent advancements for triangulations include new constructions Speleers (2013); Lyche et al. (2022); Grošelj and Speleers (2023, 2017); Lyche et al. (2025). For T-meshes, notable constructions include T-splines Sederberg et al. (2003, 2004); Li et al. (2012), HB-splines Forsey and Bartels (1988); Vuong et al. (2011), LR-splines Dokken et al. (2013), and PHT-splines Deng et al. (2008). However, except for PHT-splines, these typically do not form a complete basis for the polynomial spline space over the corresponding T-mesh Buffa et al. (2010); Mokriš et al. (2014); Giannelli and Jüttler (2013); Dokken et al. (2013). It must be emphasized that in non-basis function construction, the space spanned by linearly independent spline functions may not reproduce polynomials Sederberg et al. (2003, 2004), leading optimal convergence cannot be guaranteed.
@@
This paper extends the framework of Zhong et al. (2025) to polynomial spline spaces over arbitrary planar polygonal partitions. We observe that any polygonal partition can be transformed into a cross-cut or quasi-cross-cut partition, which possess stable dimensions and well-defined basis function structures Wang (2013); Manni (1992). By applying the Extended Edge Elimination Condition (EEE condition), we derive a basis for the spline space over the original partition. To address the complexity of the EEE condition due to excessive extensions, we propose a simplified framework based on dimensional stability. This framework also facilitates the analysis of dimensional instability.
@@
This study advances basis construction for spline spaces over arbitrary polygonal partitions, challenging the notion that basis functions cannot be constructed on dimensional unstable meshes. The key contributions are:
@@ 2 Preliminaries
Lemma 2.1.
Chui and Wang (1983a) Let $\Delta_{c}$ be a cross-cut partition of $\Omega$ with $L$ cross-cuts and $V$ interior vertices $A_{1},\ldots,A_{V}$ in $\Omega$ such that $N_{i}$ cross-cuts intersect at $A_{i}$ for $i=1,2,\ldots,V$. Then the dimension of spline space over $\Delta_{c}$ is
@@
(2.1) $k_{d}^{\mu}(N_{i})=\sum\limits_{j=1}^{d-\mu}\left(N_{i}(d-\mu-j+1)-(d-j+2)\right)_{+}.$
Lemma 2.2.
Chui and Wang (1983a); Manni (1992) Let $\Delta_{qc}$ be a quasi-cross-cut partition of $\Omega$ with $L_{1}$ cross-cuts, $L_{2}$ rays and $V$ interior vertices $A_{1},\ldots,A_{V}$ in $\Omega$ such that $N_{i}$ cross-cuts and rays intersect at $A_{i}$ for $i=1,2,\ldots,V$. Then the dimension of spline space over $\Delta_{qc}$ is
@@
Lemma 2.3.
Wang (2013) For any polygonal partition $\Delta$,let $\phi_{i},\phi_{j}$ are two adjcent cells of $\Delta$, $s(x,y)|_{\phi_{i}}=s_{i}(x,y),s(x,y)|_{\phi_{j}}=s_{j}(x,y)$, and the equation of the common edge of $\phi_{i}$ and $\phi_{j}$ is $L_{ij}(x,y)=0$. Then
@@ 3 Framework for Basis Function Construction
In this section, we introduce the concept of a Basis Construction-Suitable Partition for spline spaces, enabling direct basis construction without edge extension or elimination. We apply the Extended Edge Elimination (EEE) conditions to this polygonal partition, establishing a framework for constructing basis functions for the spline space. To clarify the construction process, we first consider the one-dimensional case before addressing the two-dimensional case.
@@ 3.1 Motivation: One-Dimensional Case
Definition 3.1.
Lyche and Morken (2008) A knot vector $\Delta_{1}$ is a refinement of $\Delta_{2}$ if every real number in $\Delta_{2}$ appears in $\Delta_{1}$.
@@
This study aims to construct a basis for the spline space $S_{d}^{\mu}(\Delta_{2})$ from $S_{d}^{\mu}(\Delta_{1})$, addressing the following problem:
(3.2) $\textbf{Construct a basis for }S_{d}^{\mu}(\Delta_{2})\textbf{ from }S_{d}^{\mu}(\Delta_{1}).$
A key observation is that the primary distinction between $S_{d}^{\mu}(\Delta_{1})$ and $S_{d}^{\mu}(\Delta_{2})$ lies in the smoothness condition: in $S_{d}^{\mu}(\Delta_{2})$, the knots in $\Delta_{1}\setminus\Delta_{2}$ are $C^{\infty}$. This leads to the following lemma for $S_{d}^{\mu}(\Delta_{2})$.
Lemma 3.4.
@@
(3.3) $\left.\frac{\partial^{r}_{+}}{\partial x^{r}}s(x)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}=\left.\frac{\partial^{r}_{-}}{\partial x^{r}}s(x)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}\quad r=\mu+1,\ldots,d.$
@@
(3.4) $\left.\sum c_{i}\left(\frac{\partial^{r}_{+}}{\partial x^{r}}B_{i,\Delta_{1}}(x)-\frac{\partial^{r}_{-}}{\partial x^{r}}B_{i,\Delta_{1}}(x)\right)\right|_{x\in\Delta_{1}\setminus\Delta_{2}}=0\quad r=\mu+1,\ldots,d,$
@@
(3.5) $\left\{\sum c_{i,j}B_{i,\Delta_{1}}(x)\right\}_{j=1}^{h}$
@@ 3.2 Two-Dimensional Case
In Zhong et al. (2025), a basis construction method for spline spaces over diagonalizable T-meshes is introduced, followed by the use of EEE conditions to construct bases for spline spaces over arbitrary T-meshes. This section generalizes the basis construction for diagonalizable T-meshes, introducing the concept of Basis Construction-Suitable Partition. We then clarify the relationship and distinctions between Dimensional-Stable Partitions and Basis Construction-Suitable Partitions. Finally, we extend the EEE conditions to arbitrary polygonal partitions, enabling basis construction for spline spaces over such partitions. Before introducing the main result, we examine the extended polygonal partition, denoted $\mathrm{ext}_{s}(\Delta)$, as follows.
Definition 3.8.
@@
$\mathrm{ext}_{0}({\Delta}):={\Delta}\subseteq\mathrm{ext}_{1}({\Delta})\subseteq\mathrm{ext}_{2}({\Delta})\subseteq\cdots\subseteq\mathrm{ext}_{s}({\Delta}):={\Delta}_{c}.$
@@
$S_{d}^{\mu}\left(\mathrm{ext}_{0}({\Delta})\right)\subseteq S_{d}^{\mu}\left(\mathrm{ext}_{1}({\Delta})\right)\subseteq\cdots\subseteq S_{d}^{\mu}\left(\mathrm{ext}_{s}({\Delta})\right).$
where $\subseteq$’ denotes a subspace relation.
We now define the Basis Construction-Suitable Partition (or Basis Construction-Suitable Mesh), which enables direct construction of basis functions for the spline space $S_{d}^{\mu}({\Delta})$ and serves as a core component in the basis construction framework.
Definition 3.10.
A polygonal partition $\Delta$ is termed a Basis Construction-Suitable Partition if:
@@
By Definitions 3.10 and 3.11, every Basis Construction-Suitable Partition is a Dimensional-Stable Partition, but the converse is not true. The Basis Construction-Suitable Partition depends on the basis construction method. A Dimensional-Stable Partition may form a Basis Construction-Suitable Partition under one construction method but not under another. Whatever, the dimensional stability condition is necessary.
Example 3.12.
@@
Zhong et al. (2025) Consider the T-mesh $\Delta$} with a T-cycle, as shown in the left subfigure of Figure 1. For the spline space $S_{2,2}^{1,1}({\Delta})$, the dimension is stable. We construct basis functions for $S_{2,2}^{1,1}({\Delta})$ using two methods:
@@
  • U-spline method Thomas et al. (2022): Basis functions are constructed by forming linear systems from smoothness constraints on the Bézier coordinates of each cell and solving these systems.
@@
    For the T-mesh $\Delta$} depicted in the left subfigure of Figure 1, this method directly yields a basis for $S_{2,2}^{1,1}({\Delta})$, confirming that $\Delta$} is a Basis Construction-Suitable Partition.
@@
  • PT-spline method Zhong et al. (2025): Basis functions are constructed by extending edges in the T-mesh to form an extended T-mesh $\mathrm{ext}_{s}({\Delta})$ that supports local tensor product B-splines, satisfying the dimension formula for $S_{\boldsymbol{d}}^{\boldsymbol{\mu}}(\mathrm{ext}_{s}({\Delta}))$. The Extended Edge Elimination Condition is then applied to obtain a basis for $S_{\boldsymbol{d}}^{\boldsymbol{\mu}}(\mathrm{ext}_{s}({\Delta}))$.
@@
    For the T-mesh $\Delta$} depicted in the left subfigure of Figure 1, edge extensions are necessary to satisfy the PT-spline construction conditions, as depicted in the right subfigure of Figure 1. Consequently, the extended T-mesh $\mathrm{ext}_{2}({\Delta})$ forms a Basis Construction-Suitable Partition. See Example 11 in Zhong et al. (2025) for details.
Dimensional-Stable PartitionBasis Construction-Suitable Partition
@@
Figure 1: Comparison of Dimensional-Stable Partition and Basis Construction-Suitable Partition
Definition 3.13.
@@
For an arbitrary polygonal partition $\Delta$, let $\mathcal{M}$ denote a basis construction method. The pair $(\Delta,\mathcal{M})$ is compatible if $\Delta$ is a Basis Construction-Suitable partition under $\mathcal{M}$.
@@
Many partitions studied for constructing spline basis functions are Basis Construction-Suitable Partitions, enabling direct basis construction. In the T-mesh category, these include hierarchical meshes with mild restrictions for HB-splines Mokriš et al. (2014), AS T-meshes for AS T-splines Li et al. (2012); Li and Scott (2014), restricted LR-meshes for LR-splines Patrizi et al. (2020), and hierarchical T-meshes for PHT-splines Deng et al. (2008). In the triangulation category, examples include Clough-Tocher (CT) splits Lai and Schumaker (2007); Sablonnière (1987) for $S_{d}^{2}(\Delta_{\mathrm{CT}})$ with $d\geq 7$ and Powell-Sabin (PS) 6 and 12 splits Powell and Sabin (1977); Alfeld and Schumaker (2002); Schumaker and Sorokina (2006); Lai and Schumaker (2007); Speleers (2013) for $S_{d}^{2}(\Delta_{\mathrm{PS}})$ with $d\geq 5$ etc.
Example 3.14.
@@
The pair $(\Delta,\mathcal{M}_{\text{HB}})$ is not compatible because $\Delta$ does not form a completeness hierarchical B-spline mesh Mokriš et al. (2014). For $S_{2}^{1}(\Delta)$, $(\Delta,\mathcal{M}_{\text{PHT}})$ is not compatible as the smoothness and degree do not satisfy PHT-spline requirements, which mandate smoothness less than half the degree. For $S_{3}^{1}(\Delta)$, $(\Delta,\mathcal{M}_{\text{PHT}})$ is compatible since $\Delta$ is a hierarchical T-mesh and the smoothness meets PHT-spline requirements.
@@
In fact, cross-cut and quasi-cross-cut partitions are both Basis Construction-Suitable Partitions for the construction method mentioned in Wang (2013). The cross-cut partition is the maximal Basis Construction-Suitable Partition for a given polygonal partition $\Delta$.
@@
Definition 3.15.
Given a polygonal partition $\Delta$} and extended Basis Construction-Suitable Partition $\mathrm{ext}_{s}({\Delta})$, as defined in Definition 3.10, assume that $\{N_{1},N_{2},\ldots,N_{n}\}$ forms a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}({\Delta}))$. For an extended edge $l$ in $\mathrm{ext}_{s}({\Delta})$, let $\phi_{1}$ and $\phi_{2}$ be two adjacent cells intersecting at $l$. The condition
(3.6) $\sum_{i=1}^{n}c_{i}N_{i}\in C^{d}(\overline{\phi_{1}\cup\phi_{2}})$
@@
(3.7) $\sum_{i=1}^{n}c_{i}\cdot\frac{\partial^{r}}{\partial\boldsymbol{n}^{r}}\left(N_{i}|_{\phi_{1}}-N_{i}|_{\phi_{2}}\right)\Big{|}_{l}\equiv 0,\quad r=\mu+1,\ldots,d,$
@@
Proof 3.17.
Since $N_{i}\in S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$, the EEE condition (3.6) implies
@@
Definition 3.18.
Given a polygonal partition $\Delta$ and extended Basis Construction-Suitable partition $\mathrm{ext}_{s}(\Delta)$, let $\{N_{1},N_{2},\ldots,N_{n}\}$ be a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$. For the set of extended edges $\{l_{1},l_{2},\ldots,l_{s}\}$, the EEE conditions for each $l_{i}$, $i=1,\ldots,s$, form the linear system
(3.8) $M\boldsymbol{c}=0,$
@@
Theorem 3.19.
Given a polygonal partition $\Delta$ and its extended Basis Construction-Suitable partition $\mathrm{ext}_{s}(\Delta)$, let $\{N_{1},N_{2},\ldots,N_{n}\}$ be a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$. If the homogeneous linear system $M\boldsymbol{c}=0$ from (3.8) has a basis of solutions $\{\tilde{c}_{j}\}_{j=1}^{h}$, where $\tilde{c}_{j}=(\tilde{c}_{1,j},\tilde{c}_{2,j},\ldots,\tilde{c}_{n,j})$, then the functions
@@
Proof 3.20.
Let $\{l_{1},l_{2},\ldots,l_{s}\}$ denote the set of all extended edges. The proof proceeds in two steps:
@@
    For any $s\in S_{d}^{\mu}(\Delta)$, Lemma 3.9 implies $s\in S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$, allowing the representation
@@
    Since $s\in S_{d}^{\mu}(\Delta)$, for each virtual edge $l_{i}$, $i=1,2,\ldots,s$, it follows $\sum_{i=1}^{n}s_{i}N_{i}\in C^{\infty}\left(\overline{\phi_{1}^{{(i)}}\cup\phi_{2}^{{(i)}}}\right)$ where $\phi_{1}^{{(i)}},\phi_{2}^{{(i)}}$ are two adjacent cells of $l_{i}$. Consequently, in $\Delta$, the following conditions hold: